One gallon equals 4 quarts. Write a rule that expresses the number of gallons g of a liquid as a function of the number of quarts q. Use the rule to find the number of gallons in 30 quarts.
Answer by ad_alta(240)

1 gallon=4 quarts, so the function G(x)=Q/4 takes in a number in "quarts" and outputs an equivalent number of "gallons." G(30)=30/4=7.5 gallons.